OpenStudy (kamille):

hey,I have system of equation to solve, can anyone have a look? \[\left(\begin{matrix}\frac{ x+2x+y }{ 2 }=2(\sqrt{6} +4\sqrt{3}) \\ x ^{2}+y ^{2}=(2x)^{2}\end{matrix}\right)\]

OpenStudy (anonymous):

You can solve second equation a bit more.. :)

OpenStudy (anonymous):

\(y^2 = 3x^2\)

OpenStudy (kamille):

so firstly: \[\frac{ 3x+y }{ 2 }=2(\sqrt{6}+4\sqrt{3})|*2\] \[3x+y=4(\sqrt{6}+4\sqrt{3})\] \[3x+y=4\sqrt{6}+16\sqrt{3}\]

OpenStudy (kamille):

then I can say that y=.. \[y=4\sqrt{6}+16\sqrt{3}-3x\]

OpenStudy (anonymous):

"Squaring it" is not solving your problem?

OpenStudy (kamille):

I dont know how to 'square' when there is three numbers:/

OpenStudy (kamille):

I only know the rule (a+b)^2=a^2+2ab+c, but I dunno what to do,when it is (a+b+c)^2

OpenStudy (kamille):

@waterineyes

OpenStudy (kamille):

@hartnn can you have a look?

hartnn (hartnn):

is that x+2x+y ?? thats just 3x+y

hartnn (hartnn):

ok, i see u got that

OpenStudy (kamille):

yeah, the original problem was a geometric one,I just need to solve this equation in oder to find the answer I found that (a+b+c)²=a²+b²+c²+2ab+2ac+2bc

hartnn (hartnn):

so, (a+b+c)^2 = a^2+b^2 +c^2 + 2ab+2bc+2ac

OpenStudy (kamille):

oh gosh, lots of work

hartnn (hartnn):

you sure, what you posted is original question ?

OpenStudy (kamille):

well, the orignal question is hard to translate, but I will try... It is the outcome of the original question

hartnn (hartnn):

thought so, you sure this outcome is correct ?

OpenStudy (kamille):

The half perimeter (p) of a right triangle is \[p=2(\sqrt{6}+4\sqrt{3)}\] One of the angle is 60 degrees (so another one is 30 degrees, because it is a right triangle)). I need to find an area of this triangle. So what I did: since I know, that one of the angle is 30 degrees,it means that the side in front of this angle is 1/2 the lenght of hypotenuse. If we say that the side in front of 30 degrees is x, then hypotenuse is 2x and another side is y (because we don'tknow anything about it) so we get the equation: \[\frac{ x+2x+y }{ 2 }=2(\sqrt{6}+4\sqrt{3})\] (it is because p is half of Perimeter) another equation is from pythagoras theorem: \[x ^{2}+y ^{2}=(2x)^{2}\]

OpenStudy (kamille):

I dunno,maybe there is an easier way @hartnn

hartnn (hartnn):

so you know side opposite to 30 degree angle is hypo/2 but you don't know side opposite to 60 degree angle is sqrt 3/ 2 times hypo ?

OpenStudy (kamille):

no,I didnt know this

OpenStudy (kamille):

I dont think they teach this at school, where I am from tho:/

hartnn (hartnn):

ok, so now you know y

OpenStudy (kamille):

so basically I dont even need the system of equation?

hartnn (hartnn):

not actually....you get the same using pythagoras theorem

OpenStudy (kamille):

I am confused, can you show me what I need to do?

hartnn (hartnn):

area is xy/2 perimeter is x+y+2x let me think of a shorter way if any

hartnn (hartnn):

instead of squaring 1st equation take square root from 2nd equation

hartnn (hartnn):

y = sqrt 3 x

hartnn (hartnn):

plug this in 1st equation much easier than squaring

OpenStudy (kamille):

oh, so y=\[y=x \sqrt{3}\]

hartnn (hartnn):

yes

OpenStudy (kamille):

and now put that in the first equation?

hartnn (hartnn):

yup

OpenStudy (kamille):

\[x \sqrt{3}=4\sqrt{16}+16\sqrt{3}+3x\] no idea how to solve this, tho

hartnn (hartnn):

its a linear equation!

OpenStudy (kamille):

but I cant sum up any of those together, can I?

hartnn (hartnn):

4 sqrt 16 or 4 sqrt 6 ??

OpenStudy (kamille):

6,sorry

OpenStudy (kamille):

ok I think I nailed it

hartnn (hartnn):

ok \(x \sqrt 3 +3x = 4 \sqrt 6 + 16 \sqrt 3 \) yes ? \(x (3+\sqrt 3) = 4 \sqrt 6 + 16 \sqrt 3\)

OpenStudy (kamille):

\[x \sqrt{3}-3x=4\sqrt{6}+16\sqrt{3}\]

OpenStudy (kamille):

yes and now you need to divide by (3+sqert 3),right?

hartnn (hartnn):

\(x \sqrt 3 (1+\sqrt 3) = 4 \sqrt 3 (\sqrt 2 + 4 \sqrt 3)\)

OpenStudy (kamille):

oh wait how did you get it

hartnn (hartnn):

why is there a -3x ?

OpenStudy (kamille):

I would have done somethinglike that: \[x(\sqrt{3}+3=4(\sqrt{6}+4\sqrt{3})|:(\sqrt{3}+3)\]

hartnn (hartnn):